If you wish to withdraw Rs. 50000, Rs 80,000, Rs. 110000 and Rs. 140000 at the end of 2nd, 3rd, 4th and 5th year from now from a savings accounts which earns 8% interest rate compounded annually, the amount you should deposit now is


1
Expert's answer
2021-02-25T05:07:37-0500

"A=P(1+r\/n)^(nt)"

To withdraw Rs. 50,000 at the end of 2nd year

"50,000=P(1+0.08)^2"

"50,000=1.1664P"

"P=50,000\u00f71.1664"

"P=Rs. 42,866.94"


To withdraw Rs. 80,000 at the end of 3 years

"80,000=P(1+0.08)^3"

"P=80,000\u00f7(1.08)^3"

"P=Rs.63,506.58"


To withdraw Rs. 110,000 at the end of 4 years

"P=110,000\u00f7(1.08)^4"

"P=Rs.80,853.28"


To withdraw Rs. 140,000 at the end of 5 yearsyears

"P=140,000\u00f7(1.08)^5"

"P=Rs.95, 281.65"